OpenStudy (anonymous):

Please Help!! Part A: Factor 3x2y2 - 2xy2 - 8y2. Show your work. (3 points) Part B: Factor x2 + 10x + 25. Show your work. (3 points) Part C: Factor x2 - 36. Show your work. (2 points) Part D: Do the polynomials in parts A, B, and C have a common factor? Explain your answer. (2 points)

OpenStudy (anonymous):

I already got part A

OpenStudy (anonymous):

I only need part b, c and d

OpenStudy (anonymous):

Ahem well i'll contribute to B: \[x^2+10x+25=(x+5)(x+5)=(x+5)^2\]

OpenStudy (anonymous):

Oh and i'll do C as well: anything in the form of \((x^2-a^2)\) where a is an integer is factored as \((x+a)(x-a)\) \[x^2-36=x^2-(6)^2=(x+6)(x-6)\]

OpenStudy (lanhikari22):

that's \[x^2\] ? i thought it was a 2x v.v

OpenStudy (anonymous):

Thats it for b and c?

OpenStudy (anonymous):

@KeithAfasCalcLover

OpenStudy (anonymous):

I need part D, Anyone?

OpenStudy (anonymous):

Please I need help..

OpenStudy (anonymous):

Yeah thats all there is! Umm...no there is no common factor in all of them...

OpenStudy (anonymous):

ok thank you so much! @KeithAfasCalcLover would you help me with one more?

OpenStudy (anonymous):

Linda is studying the sale of a particular brand of cereals from the year 1993 to 2004. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t2 + 5t + 52 Part A: What does the y-intercept of the graph of the function represent? (4 points) Part B: What is the reasonable domain of the graph of the function? (3 points) Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work. (3 points)

OpenStudy (anonymous):

So we know "to find the y-intercept, set t to zero" Therefore, what they are asking is for part a is when t=0 (or at zero seconds) what is the sale of the cereal. And in this case, its 52. So when they started this graph, the cereal was selling 52,000,000 dollars. Now b would be what is the domain.  where the parabola equals zero would be your domain. So find between what x values \(t^2+5t+52\)=0 and thats your answer. C uses the formula: The average rate of change, m, of \(f(x)\) from x=a to b is: \[m=\frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

So what is the answer for part A?

OpenStudy (anonymous):

I didnt really understand what you wrote... @KeithAfasCalcLover

OpenStudy (anonymous):

I understood a little tho

OpenStudy (anonymous):

Offline... Thanx anyway